# KLAN: Kuaishou Landing-page Adaptive Navigator

### v1

@@ 6. Conclusion and Future Work
In this paper, we formally introduce Personalized Landing Page Modeling (PLPM) as a critical yet underexplored task in recommender systems. To address PLPM, we propose KLAN, a hierarchical framework that real-time integrates inter-day static page preferences with intra-day dynamic interests through adaptive modulation. Extensive online experiments on Kuaishou demonstrate KLAN’s effectiveness and scalability. Last but not least, our model is now deployed online with full users at Kuaishou platform, serving for hundreds of millions of users every day. Additionally, we will provide a well-established benchmark for the PLPM problem to encourage future innovations.
@@ A.2. Feature Embedding and Selection
[…]
@@ A.3. KLAN-ISP Theoretical Analysis
[…]
An example of the gradient conflict.
@@ A.4. KLAN-IIT Design Details
(26) $V_{t}=\frac{24*\text{The total usage time of the $t$-th hour}}{\text{The total usage time of the whole day}}$
@@
where $\alpha$ is the base coefficient. We constrain $\alpha_{t}$ within bounds of $[0.1\alpha,5\alpha]$ to ensure stability.
@@ A.7. Notations and Corresponding Descriptions
$N$ Number of all instances in $\mathcal{Z}$

### v2

@@ 6. Conclusion and Future Work
In this paper, we formally introduce Personalized Landing Page Modeling (PLPM) as a critical yet underexplored task in recommender systems. To address it, we propose KLAN, a hierarchical framework that integrates inter-day static page preferences with intra-day dynamic interests through adaptive modulation in real-time. Extensive online experiments on Kuaishou demonstrate KLAN’s effectiveness and scalability. Moreover, our model is now deployed online with full users at Kuaishou platform, serving for hundreds of millions of users every day. Finally, we will also release a well-established benchmark to encourage future innovations in the PLPM scenario.
@@ A.2. Feature Embedding and Selection
$\left\{\begin{aligned} \mathbf{w}_{i}^{k}&=Softmax\left(\mathbf{W}^{k}\cdot\mathbf{e}_{ti}+\mathbf{b}^{k}\right)\\ \mathbf{e}_{xi}^{k}&=\sum_{j=1}^{f_{x}}\left(\mathbf{w}_{i}^{k}(j)\cdot\mathbf{e}_{xi,j}\right)\end{aligned}\right$.
@@ A.3. KLAN-ISP Theoretical Analysis
Refer to caption An example of the gradient conflict.
@@ A.4. KLAN-IIT Design Details
[…]
@@
where $\alpha$ is the base coefficient. We constrain $\alpha_{t}$ within bounds of $0.1\alpha,5\alpha]$ to ensure stability.
@@ A.7. Notations and Corresponding Descriptions
NN Number of all instances in $\mathcal{Z}$